# Feasibility Study of Measuring $\Lambda^{0}\to n\pi^{0}$ Using a High-Granularity Zero-Degree Calorimeter at the Future Electron-Ion Collider

### v1

@@ 1 Introduction
In contrast, the prospects of measuring $\Lambda^{0}\to n\pi^{0}$ (branching ratio 36% [9]) with the ZDC has not been studied with detailed simulations. For instance, the EIC YR [4] only presents kinematic distributions for the decay products of $\Lambda^{0}\to n\pi^{0}$ without detector simulations, from which only limited information can be inferred. Similarly, Ref. [13] presents a similar study for the EicC. To date, no prior study has demonstrated feasibility taking into account detector response or clustering algorithms and particle identification algorithms
@@ 2.1 Kinematics of $\Lambda^{0}$ decay
The $\Lambda^{0}$ then decays isotropically in its rest frame into a neutron and a $\pi^{0}$, which promptly decays isotropically in its rest frame into two photons. Consequently, the neutron carries between 75% and 94% of the energy of the $\Lambda^{0}$ in the lab frame, regardless of the momentum of the $\Lambda^{0}$. This is because the neutron is much more massive than the $\pi^{0}$, and in the $\Lambda^{0}$ center-of-mass frame, the momenta of both of these two particles is around 100 MeV, which is relatively small compared to the mass of the neutron.
@@ 2.2 Detector Response
Following our previous studies [14, 20], hits in the ZDC are reconstructed using a 15-bit ADC with no noise and a dynamic range of 800 MeV. These hits are included in the subsequent event reconstruction only if they exceed an energy threshold of 0.5 MIPs [^4] One MIP is the most-probable hit energy deposited by a minimum-ionizing particle, which was determined using muon simulations to be 0.47 MeV., and the the hit time is less than 270 ns555that is, the amount of time for a particle moving at the speed of light to reach the ZDC front face from the interaction point, plus an additional 150 ns.. These cuts were inspired by those used in CALICE prototypes [17, 18] and the anticipated capabilities of the EIC ASIC for SiPM readout. Hits passing this selection are simply referred as hits in the rest of this work.
@@ 3.1 Geometric Acceptance
The geometrical acceptance for $\Lambda^{0}\rightarrow n\pi^{0}$ is defined as the probability that the $\Lambda^{0}$ decays before reaching the front face of the ZDC and that all three daughter particles strike the ZDC within its fiducial region [^6] This fiducial region is determined by the aperture of the magnets and the design of the beam-pipe exit window before the ZDC., defined by a radius of 14.3 cm.
@@ 3.2 Energy reconstruction and Clustering
We selected events with two photon clusters and at least one other cluster [^7] Hadronic showers in a high-granularity calorimeter like the ZDC [14] are expected to yield multiple clusters when using the topocluster algorithm.. The energies of the photons were then the energies of their respective clusters. All other clusters are then associated with the neutron. The first step to obtaining the neutron’s energy is adding up all the non-photon clusters. A correction factor is applied to the reconstructed neutron energy to account for the non-compensated nature of the ZDC  [^8] This is a first-order correction, and more refined methods, such as software compensation [19] that weights each topocluster differently, could be applied to improve approach in future work.:
@@ 3.3 Displaced Vertex and Mass Reconstruction
where $\vec{p}_{i}$ is the momentum of the particle, $E_{i}$ is the energy, $m_{i}$ is the mass of the particle from the PDG (0 for photons, and $\approx 941$ MeV for neutrons [9]), $\vec{x}^{\,\,i}_{\rm cl}$ is the position of the cluster [^10] The position the highest energy cluster associated with the neutron., and $\vec{x}_{\rm vtx}$ is the decay position in this iteration. We then obtain from these momenta the momentum vector of the $\Lambda^{0}$ as the sum of those of the three particles,
@@ 4.1 Lambda identification
At 100 GeV and $p_{T}$=0 GeV, the combined acceptance and efficiency is 2%$\times$9or 0.018% for conventional methods and 2%$\times$40%, or 0.8% for the GNN method. These values increase to around 35%$\times$5$\approx$1.7for conventional and 35%$\times$70$\approx$24at 250 GeV. Assuming an integrated luminosity of 10 fb-1, a cross-section of roughly 100 nb [^12] Which is consistent with estimates of the $\Lambda^{0}$ production cross section from the DEMPgen event generator [39] for the EIC at the 18 $\times$ 275 GeV setting., and that the combined acceptance $\times$ efficiency averaged over the $\Lambda^{0}$ energy and $p_{T}$ range is $\approx$10%, then we would expect a sample size of around 100k events.
@@ 4.4 Transverse momentum resolution
However, due to the spread of the proton beam, $\sigma_{\rm beam}=\mathscr{O})$(100 $\mu$rad), the actual $p_{T}$ resolution is worse than the ZDC resolution. With beam effects, the expression for the $p_{T}$ resolution becomes:
@@ 5 Summary
We have conducted a feasibility study for reconstructing high-energy $\Lambda^{0}$ baryons at the EIC through their neutral decay channel, $\Lambda^{0}\to n\pi^{0}$, using a high-granularity Zero Degree Calorimete [14]. The main challenge with this channel is estimating the $\Lambda^{0}$ displaced vertex, which, at the relevant EIC energies, is expected to range from meters to tens of meters—potentially the longest distances ever measured for displaced vertices in any experiment. We addressed this challenge using an iterative method that explicitly estimates the displaced vertex, as well as an AI-based approach that resolves it implicitly.
@@
Furthermore, we demonstrated that the $\Lambda^{0}\to n\pi^{0}$ decay channel is feasible using only the SiPM-on-tile ZDC design [14], contradicting the EIC Yellow Report’s claim that measuring this channel “requires a means for EM Calorimetry before the ZDC, in order to distinguish the neutron and the two photons coming from $\pi^{0}$ decay” [4].

### v2

@@ 1 Introduction
In contrast, the prospects of measuring $\Lambda^{0}\to n\pi^{0}$ (branching ratio 36% [9]) with the ZDC has not been studied with detailed simulations. For instance, the EIC YR [4] only presents kinematic distributions for the decay products of $\Lambda^{0}\to n\pi^{0}$ without detector simulations, from which only limited information can be inferred. Arrington et al. suggested from this kinematics study the possibility of measuring this decay channel combining the measurement of the neutron in the ZDC with the measurement of the photons from the $\pi^{0}$ decay in an electromagnetic calorimeter in the B0 region [7]. Similarly, Ref. [13] presents a similar study for the EicC. To date, no prior study has demonstrated feasibility taking into account detector response or clustering algorithms and particle identification algorithms.
@@ 2.1 Kinematics of $\Lambda^{0}$ decay
In the simulation, the $\Lambda^{0}$ then decays isotropically in its rest frame into a neutron and a $\pi^{0}$, which promptly decays isotropically in its rest frame into two photons. Consequently, the neutron carries between 75% and 94% of the energy of the $\Lambda^{0}$ in the lab frame, regardless of the momentum of the $\Lambda^{0}$. This is because the neutron is much more massive than the $\pi^{0}$, and in the $\Lambda^{0}$ center-of-mass frame, the momenta of both of these two particles is around 100 MeV, which is relatively small compared to the mass of the neutron.
@@ 2.2 Detector Response
Following our previous studies [14, 20], hits in the ZDC are reconstructed using a 15-bit ADC with no noise and a dynamic range of 800 MeV. These hits are included in the subsequent event reconstruction only if they exceed an energy threshold of 0.5 MIPs [^4] One MIP is the most-probable hit energy deposited by a minimum-ionizing particle, which was determined using muon simulations to be 0.47 MeV., and the hit time is less than 270 ns555that is, the amount of time for a particle moving at the speed of light to reach the ZDC front face from the interaction point, plus an additional 150 ns.. These cuts were inspired by those used in CALICE prototypes [17, 18] and the anticipated capabilities of the EIC ASIC for SiPM readout. Hits passing this selection are simply referred as hits in the rest of this work.
@@ 3.1 Geometric Acceptance
The geometrical acceptance for $\Lambda^{0}\rightarrow n\pi^{0}$ is defined as the probability that the $\Lambda^{0}$ decays before reaching the front face of the ZDC and that all three daughter particles strike the ZDC within its fiducial region [^6] This fiducial region is determined by the aperture of the magnets and the design of the beam-pipe exit window before the ZDC., defined by a radius of 14.3 cm. For comparison, the transverse size of the detector is $\approx 60\times 60$ cm2.
@@ 3.2 Energy reconstruction and Clustering
We selected events with two photon clusters and at least one other cluster [^7] Hadronic showers in a high-granularity calorimeter like the ZDC [14] are expected to yield multiple clusters when using the topocluster algorithm.. The energy of each reconstructed photon, prior to corrections, is the energy of its associated cluster. All clusters not associated with a photon were then associated with the neutron. The first step to obtaining the neutron’s energy is adding up the energies of all the non-photon clusters. A correction factor is applied to the reconstructed neutron energy to account for the non-compensated nature of the ZDC  [^8] This is a first-order correction, and more refined methods, such as software compensation [19] that weights each topocluster differently, could be applied to improve approach in future work.:
@@ 3.3 Displaced Vertex and Mass Reconstruction
where $\vec{p}_{i}$ is the momentum of the particle, $E_{i}$ is the energy, $m_{i}$ is the mass of the particle from the PDG (0 for photons, and $\approx 940$ MeV for neutrons [9]), $\vec{x}^{\,\,i}_{\rm cl}$ is the position of the cluster [^10] For neutrons, which can have more than one cluster, this is the of position the highest-energy cluster associated with the neutron., and $\vec{x}_{\rm vtx}$ is the decay position in this iteration. We then obtain from these momenta the momentum vector of the $\Lambda^{0}$ as the sum of those of the three particles,
@@ 4.1 Lambda identification
At 100 GeV and $p_{T}$=0 GeV, the combined acceptance and efficiency is 2%$\times$9or 0.018% for conventional methods and 2%$\times$40%, or 0.8% for the GNN method. These values increase to around 35%$\times$5$\approx$1.7for conventional and 35%$\times$70$\approx$24at 250 GeV. Assuming an integrated luminosity of 10 fb-1, a cross-section of roughly 100 nb [^12] Which is consistent with estimates of the $\Lambda^{0}$ production cross section from the DEMPgen event generator [39] for the EIC at the 18 $\times$ 275 GeV setting., a $\Lambda^{0}\rightarrow n\pi^{0}$ branching ratio of 36% [9], and that the combined acceptance $\times$ efficiency averaged over the $\Lambda^{0}$ energy and $p_{T}$ range is $\approx$10%, then we would expect a sample size of around 36k events.
@@ 4.4 Transverse momentum resolution
However, due to the spread of the proton beam, $\sigma_{\rm beam}=\mathscr{O}$(100 $\mu$rad), the actual $p_{T}$ resolution is worse than the ZDC resolution. With beam effects, the expression for the $p_{T}$ resolution becomes:
@@ 5 Summary
We have conducted a feasibility study for reconstructing high-energy $\Lambda^{0}$ baryons at the EIC through their neutral decay channel, $\Lambda^{0}\to n\pi^{0}$, using a high-granularity Zero Degree Calorimeter [14]. The main challenge with this channel is estimating the $\Lambda^{0}$ displaced vertex, which, at the relevant EIC energies, is expected to range from meters to tens of meters—potentially the longest distances ever measured for displaced vertices in any experiment. We addressed this challenge using an iterative method that explicitly estimates the displaced vertex, as well as an AI-based approach that resolves it implicitly.
@@
Furthermore, we demonstrated that the $\Lambda^{0}\to n\pi^{0}$ decay channel is feasible using the SiPM-on-tile ZDC design [14] as a stand-alone detector which provides the dual function of hadronic and electromagnetic calorimetry at the higher proton energy.
@@
Code Availability
The code for data processing, model training, and model inference are found at https://github.com/eiccodesign/regressiononly/tree/lambda_decay. The GNN models stored as TensorFlow models and in the ONNX format are available in [40]. A stand-alone demo of the IDOLA algorithm, presented here, can be found in [41].